OpenStudy (anonymous):

Evaluate the expression when m = 4. 35 – 3m

OpenStudy (anonymous):

@aahmed123 @mastermindkakashi @shinebrightlikeadimon

OpenStudy (anonymous):

ok lets see

OpenStudy (anonymous):

Is it 1

OpenStudy (anonymous):

3 5- 3 x 4

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

not 1

OpenStudy (anonymous):

3 x 4 = 12

OpenStudy (anonymous):

35 - 3 = 32

OpenStudy (anonymous):

35 - 12 = your answer

OpenStudy (anonymous):

i mean is it 23

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

OH YEH my bad

OpenStudy (anonymous):

35 - 12 = 23

OpenStudy (anonymous):

ok next question

OpenStudy (anonymous):

new tab please

OpenStudy (anonymous):

k